A lady's purse contained 12 coins consisting of quarters and dimes only. Their total value was $1.95. How many coins of each kind were there?
Answer by mananth(16949)

let quarters be x
dimes will be 12-x
..
25x+10(12-x)= 195
25x+120-10x=195
15x=195-120
15x=75
x= 75/15
x=5 quarters
dimes = 14